z^2/8 - 7z/8 = 1 Can you please solve and show the work on how to slve this problem please
Answer by checkley77(12569)

z^2/8 - 7z/8 = 1
(z^2-7z)/8=1
z^2-7z=8
z^2-7z-8=0
(z-8)(z+1)=0
z-8=0
z=8 ans.
z+1=0
z=-1 ans.
Proofs:
8^2/8-7*8/8=1
64/8-56/8=1
(64-56)/8=1
8/8=1
1=1
-1^2/8-7*-1/8=1
1/8+7/8=1
8/8=1
1=1